OpenStudy (crashonce):

The nth term of a sequence is (5n+k) where k is a constant. If the 3rd term is -7, then the 10th term is ?

OpenStudy (crashonce):

@Mimi_x3 @perl

OpenStudy (mimi_x3):

what's \(k\)?

OpenStudy (crashonce):

constant

OpenStudy (mimi_x3):

\[(5(3)+ k) = -7 => 15 + k = -7 => k = -7 - 15\]

OpenStudy (mimi_x3):

i dont have a calcuator...can u evaluate that \(k\)

OpenStudy (mimi_x3):

ok \(k = -22\)

OpenStudy (crashonce):

wait wat do u mean k = -22

OpenStudy (mimi_x3):

\[5(10) + -22 = t\]

OpenStudy (mimi_x3):

solve for \(t\)

OpenStudy (crashonce):

28

OpenStudy (mimi_x3):

i'll trust your ansewr for that.